OpenStudy (anonymous):

jack usually mows his lawn in 6hr. mary mows the same yard in 5hr. how long will it take them to mow the yard together? i got 2 8/11?

OpenStudy (amistre64):

how much does J cut in 1 hour? how much does M cut in 1 hour? J+M = total cut in 1hour

OpenStudy (amistre64):

flip that number for time taken ...

OpenStudy (amistre64):

\[\frac{1}{6}+\frac{1}{5}=\frac{5+6}{5*6}\] flips to:\[\frac{30}{11}\]hours to complete

OpenStudy (amistre64):

your number looks good

OpenStudy (anonymous):

so would i write 2 8/11 for my answer?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

oh i see if i divide it out from 30/11 it will be that

OpenStudy (amistre64):

they tend to want the time in hrs min sec

OpenStudy (anonymous):

it will be mary plus jack= 11 hours together

OpenStudy (amistre64):

2 hours and 8/11 of another

OpenStudy (amistre64):

no abdullah; thats incorrect

OpenStudy (amistre64):

if 2 people work on the same yard they dont finish it in longer time then if they do it alone

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

well, not mathically at least lol

OpenStudy (amistre64):

60/11 = an hour divided up into 11th parts; times 8 = minutes taken in an hour

OpenStudy (amistre64):

2 hours, 43 mins, and a little over 30 secs more

OpenStudy (anonymous):

i get it